### v1

@@ Appendix A Proofs for Section 2
$\beta_{\text{TWFE}}$ $=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}]}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}]}$
$=\frac{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}\mid P=t]\mathbb{P}(P=t)}{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}\mid P=t]\mathbb{P}(P=t)}$
$=\frac{\mathbb{E}[\ddot{D}Y]}{\mathbb{E}[\ddot{D}^{2}]},$
where the second equality follows from the uniform distribution of $P$ which is independent from $(\ddot{D}_{t},Y_{t})$ for all $t\in\{1,\ldots,T\}$. The third equality follows from defining $\ddot{D}\equiv\ddot{D}_{P}$. We also note that
$\ddot{D}_{P}$ $=D_{P}-\frac{1}{T}\sum_{s=1}^{T}D_{s}-\sum_{t=1}^{T}\mathbb{E}[D_{t}]\mathbbm{1}(P=t)+\sum_{s=1}^{T}\mathbb{E}[D_{s}]\mathbb{E}[\mathbbm{1}(P=s)]$
@@
The third equality follows from $\mathbb{E}[D\mid G]=\mathbb{E}[\mathbbm{1}(G\leq P)\mid G]=\frac{1}{T}\sum_{s=1}^{T}\mathbbm{1}(G\leq s)=\frac{1}{T}\sum_{s=1}^{T}D_{s}$.
We break down the rest of this proof into four steps.
@@
$\mathbb{E}[\ddot{D}Y]=\mathbb{E}[\ddot{D}(Y(0)+D(Y(1)-Y(0)))]=\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]+\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid P,G]]$
@@
$\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]$ $=\mathbb{E}[\ddot{D}\theta(G,P)]=\mathbb{E}[\ddot{D}\ddot{\theta}(G,P)]$
@@
$\ddot{\theta}(G,P)$ $\equiv\theta(G,P)-\mathbb{E}[\theta(G,P)\mid G]-\mathbb{E}[\theta(G,P)\mid P]+\mathbb{E}[\theta(G,P)]$
@@
$\theta(g^{\prime},t^{\prime})$ $=\mathbb{E}[Y(0)\mid G=g^{\prime},P=t^{\prime}]-\mathbb{E}[Y(0)\mid G=g^{\prime}]-\mathbb{E}[Y(0)\mid P=t^{\prime}]+\mathbb{E}[Y(0)]$
$=\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g^{\prime}]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g)$
$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g).$
@@
$\theta(g^{\prime},t^{\prime})$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g)$
$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]\right)\mathbb{P}(G=g)$
@@
$\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$ $=\mathbb{E}[(D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[D(Y(1)-Y(0))\mid G,P]]$
@@
The first equality is by definition, the second by $D^{2}=D$, the third by $\mathbb{E}[D\mid G,P]=D$, and the fourth by law of total probability.
@@
$\mathbb{E}[\ddot{D}^{2}]$ $=\mathbb{E}[(D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)].$
The first line is obtained from properties of linear projections, the second from $D^{2}=D$, and the third from $D=\mathbb{P}(D=1\mid G,P)$.
We conclude the proof by noting the equivalence of integrating over the distribution of $P$ and averages over time periods, which shows the equivalence between
$\beta_{\text{TWFE}}=\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]\mathbb{P}(D=1\mid G,P)]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)]}$
and the expression in Proposition 2.4. ∎
@@
Proposition 2.4 and $\mathbb{P}(D=1\mid G,P)=D$ yielded
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]}.$
@@
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G]\mathbb{E}[Y(1)-Y(0)\mid G,D=1]]}{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G]]}.$
We now calculate the conditional expectation $\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G=g]$ for $g\in\mathcal{G}$. If $g=+\infty$, then this conditional expectation is 0 by construction, so we let $g\in\{2,\ldots,T\}$ in what follows:
@@
$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]-\frac{\mathbb{E}[F_{G}(P)\mathbbm{1}(G\leq P)\mid G=g]}{\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]}+\mathbb{E}[\mathbb{E}[\mathbbm{1}(G\leq P)\mid P]]\right)$
@@
$=\mathbb{P}(D=1\mid G=g)\mathbb{P}(D=0\mid G=g)(\mathbb{P}(D=1\mid P<g)+\mathbb{P}(D=0\mid P\geq g)).$
@@
The first equality follows from $\mathbb{E}[D\mid G=g]>0$ for $g\in\{2,\ldots,T\}$. The second follows from $D=\mathbbm{1}(G\leq P)$ and the law of iterated expectations, the third from $G\perp\!\!\!\perp P$, the fourth from definitions of conditional expectations and the law of iterated expectations, the fifth from combining terms, the seventh from the law of iterated expectations again, and the last line is obtained by the fact that $D\in\{0,1\}$. The representation in Proposition 2.5 follows. ∎
Appendix B Proofs for Section 3
Proof of Proposition 3.1.
We begin by showing the first claim of the proposition. The equation
$\mathbb{E}[(Y(1)-Y(0))W^{*}\mid W_{0}=1,X]$ $=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ (B.1)
holds since $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$, which holds by Definition 3.1. Since $\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ is assumed positive, we can divide both sides of equation (B.1) by it and obtain
$\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ $=\frac{\mathbb{E}[(Y(1)-Y(0))W^{*}\mid W_{0}=1,X]}{\mathbb{P}(W^{*}=1\mid W_{0}=1,X)}$
$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,W_{0}=1,X]$
$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,X],$
where the second equality holds by definition, and the third holds from $W^{*}$ being a subpopulation of $W_{0}$.
We now show the second claim of the proposition. Equation (3.4) can be obtained as follows:
$=\frac{\mathbb{E}[W^{*}(Y(1)-Y(0))\mid W_{0}=1]}{\mathbb{E}[W^{*}\mid W_{0}=1]}$
$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid W_{0}=1,X)\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mid W_{0}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid W_{0}=1,X)\mid W_{0}=1]}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=\mu(\underline{w}^{*},\tau_{0}).$
The first equality follows from $W^{*}$ being a subpopulation of $W_{0}$, the second from the definition of conditional expectation and $\mathbb{P}(W^{*}=1)>0$, the third from the law of iterated expectations and $W^{*}=1$ implying $W_{0}=1$, the fourth from $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$, and the fifth and sixth follow immediately. ∎
Proof of Lemma 3.1.
We verify that $W^{*}$ satisfies the two conditions in Definition 3.1. Condition 1 holds since $\mathbb{P}(W^{*}=1\mid W_{0}=0)=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid W_{0}=1)\leq\mathbb{P}(W^{\prime}=1\mid W_{0}=1)=0$, The first equality follows from $W^{*}\leq W^{\prime}$ almost surely, and the last one from $W^{\prime}\in\text{SP}(W_{0})$.
To verify condition 2, note that
$\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W_{0}=1)$ $=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W^{\prime}=1,W_{0}=1)\mathbb{P}(W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W^{\prime}=1)\mathbb{P}(W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1,W_{0}=1)\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid X,W_{0}=1).$
The first and seventh line follow from $W^{*}\leq W^{\prime}$ almost surely. The second and sixth line follow from factoring conditional probabilities. The third and fifth line follow from $W^{\prime}\leq W_{0}$ almost surely. The fourth line follows from $W^{\prime}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$ and $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W^{\prime}=1$. ∎
Proof of Theorem 3.1.
This theorem follows as a special case of the first part of Theorem 4.3 when $W^{\prime}=W_{0}$. This is because $W_{0}$ trivially regular subpopulation of $W_{0}$, and because the condition
$\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$
is equivalent to $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$. They are equivalent since $\underline{w}^{\prime}(X)=\mathbb{P}(W_{0}=1\mid X,W_{0}=1)=1$ almost surely. ∎
Proof of Theorem 3.2.
This is a special case of the second part of Theorem 4.3 where we set $W^{\prime}=W_{0}$. ∎
Proof of Proposition 3.2.
First, we suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$. Therefore, by Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{lin}}$ where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$. Therefore,
00 0 $=\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})$
$=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
for all $d\in\mathbb{R}^{d_{X}}$, which implies that $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}=\frac{\mathbb{E}[\underline{w}^{*}(X)X\mid W_{0}=1]}{\mathbb{E}[w^{*}(X)\mid W_{0}=1]}$. Letting $u(x)=\underline{w}^{*}(x)/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$, we have that

### v3

@@ Appendix E Proofs for Section 5
$\beta_{\text{TWFE}}$ $=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}]}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}]}=\frac{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}\mid P=t]\mathbb{P}(P=t)}{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}\mid P=t]\mathbb{P}(P=t)}=\frac{\mathbb{E}[\ddot{D}Y]}{\mathbb{E}[\ddot{D}^{2}]},$
where the second equality follows from the uniform distribution of $P$ which is independent from $(\ddot{D}_{t},Y_{t})$ for all $t\in\{1,\ldots,T\}$. The third equality follows from defining $\ddot{D}\coloneqq\ddot{D}_{P}$. We also note that
$\ddot{D}$ $=D_{P}-\frac{1}{T}\sum_{s=1}^{T}D_{s}-\sum_{t=1}^{T}\mathbb{E}[D_{t}]\mathbbm{1}(P=t)+\sum_{s=1}^{T}\mathbb{E}[D_{s}]\mathbb{E}[\mathbbm{1}(P=s)]$
@@
The third equality follows from $\mathbb{E}[D\mid G]=\mathbb{E}[\mathbbm{1}(G\leq P)\mid G]=\frac{1}{T}\sum_{s=1}^{T}\mathbbm{1}(G\leq s)=\frac{1}{T}\sum_{s=1}^{T}D_{s}$. We break down the rest of this proof into four steps.
@@
$\mathbb{E}[\ddot{D}Y]$ $=\mathbb{E}[\ddot{D}(Y(0)+D(Y(1)-Y(0)))]$
$=\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]+\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid G,P]].$
@@
$\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]$ $=\mathbb{E}[\ddot{D}\theta(G,P)]=\mathbb{E}[\ddot{D}\ddot{\theta}(G,P)],$
@@
$\ddot{\theta}(G,P)$ $\coloneqq\theta(G,P)-\mathbb{E}[\theta(G,P)\mid G]-\mathbb{E}[\theta(G,P)\mid P]+\mathbb{E}[\theta(G,P)]$
@@
$\ddot{\theta}(g^{\prime},t^{\prime})$ $=\mathbb{E}[Y(0)\mid G=g^{\prime},P=t^{\prime}]-\mathbb{E}[Y(0)\mid G=g^{\prime}]-\mathbb{E}[Y(0)\mid P=t^{\prime}]+\mathbb{E}[Y(0)]$
$=\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g^{\prime}]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g^{\prime}]$
$-\sum_{g\in\mathcal{G}}\left(\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g)$
$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G]\right].$
@@
$\theta(g^{\prime},t^{\prime})$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G]\right]$
$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]-\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]\right]$
@@
$\ddot{D}D$ $=(D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D$
$=(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)$ (A.1)
by $D^{2}=D=\mathbb{E}[D\mid G,P]$. Thus,
$\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$
@@
$\mathbb{E}[\ddot{D}^{2}]$ $=\mathbb{E}[\ddot{D}D]=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)].$
The first equality is obtained from properties of linear projections and the second follows from equation (A.1).
We conclude the proof by noting the equivalence of integrating over the distribution of $P$ and averages over time periods, which shows the equivalence between $\beta_{\text{TWFE}}$, the expression in Proposition 5.3, and
$\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]\cdot\mathbb{P}(D=1\mid G,P)]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot\mathbb{P}(D=1\mid G,P)]}.$
@@
Proposition 5.3 and $\mathbb{P}(D=1\mid G,P)=D$ yields
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot D\cdot\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot D]}.$
@@
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot D\mid G]\cdot\mathbb{E}[Y(1)-Y(0)\mid G,D=1]]}{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot D\mid G]]}.$
We now calculate the conditional expectation $\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot D\mid G=g]$ for $g\in\mathcal{G}$. If $g=+\infty$, then this conditional expectation is 0 by construction, so we focus on the case where $g\in\{2,\ldots,T\}$. For these derivations, we let $F_{G}(p)\coloneqq\mathbb{P}(G\leq p)$ denote the cdf of $G$ at $p$. We have that:
@@
$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]-\frac{\mathbb{E}[F_{G}(P)\mathbbm{1}(G\leq P)\mid G=g]}{\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]}\right.$
$\quad\left.+\;\mathbb{E}[\mathbb{E}[\mathbbm{1}(G\leq P)\mid P]]\right)$
@@
$=\mathbb{P}(D=1\mid G=g)\cdot\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=1\mid P<g)+\mathbb{P}(D=0\mid P\geq g)).$
@@
The first equality follows from $\mathbb{E}[D\mid G=g]>0$ for $g\in\{2,\ldots,T\}$, the second from $D=\mathbbm{1}(G\leq P)$ and the law of iterated expectations, the third from $G\perp\!\!\!\perp P$, the fourth from the law of iterated expectations, the fifth from combining terms, the sixth from the law of iterated expectations again, and the last line is obtained by the fact that $D\in\{0,1\}$. The representation in Proposition 5.4 follows. ∎
Appendix F Proofs for Appendix B
Proof of Proposition B.1.
By Theorem 3.1, $\mu(a,\tau_{0})$ has a uniform causal representation in $\mathcal{T}_{\text{all}}$ if and only if $a(x_{k})\geq 0$ for $k\in\{1,\ldots,K\}$. Therefore, it is sufficient to show the equivalence between weakly causal estimands and estimands with nonnegative weights. A similar result was shown in Proposition 4 of BBMT, but we nevertheless provide a proof here to account for the slight differences in notation.
Suppose $\mu(a,\tau_{0})$ is weakly causal. Let $\nu_{1}=(\mathbbm{1}(a(x_{1})<0),\ldots,\mathbbm{1}(a(x_{K})<0))$ and $\nu_{0}=\textbf{0}_{K}$. Trivially, $(\nu_{1},\nu_{0})\in\mathcal{M}_{\text{all}}$ and $\tau^{-}\coloneqq\nu_{1}-\nu_{0}\in\mathcal{T}_{\text{all}}$, where $\tau^{-}\geq\textbf{0}_{K}$. Since $\mu(a,\tau_{0})$ is weakly causal, $\mu(a,\tau^{-})\geq 0$ where
$\mu(a,\tau^{-})$ $=\frac{\mathbb{E}[a(X)\tau^{-}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$
$=\frac{1}{\mathbb{E}[a(X)\mid W_{0}=1]}\sum_{k=1}^{K}a(x_{k})\mathbbm{1}(a(x_{k})<0)\mathbb{P}(X=x_{k}\mid W_{0}=1)\geq 0.$
This implies $a(x_{k})\geq 0$ for all $k\in\{1,\ldots,K\}$. Thus, $\mu(a,\tau_{0})$ has a uniform causal representation in $\mathcal{T}_{\text{all}}$.
Now suppose $\mu(a,\tau_{0})$ has a uniform causal representation in $\mathcal{T}_{\text{all}}$, or that $a(x_{k})\geq 0$ for $k\in\{1,\ldots,K\}$. Then, for any $(\nu_{0},\nu_{1})\in\mathcal{M}_{\text{all}}$ such that $\tau\coloneqq\nu_{1}-\nu_{0}\geq\textbf{0}_{K}$,
$\mu(a,\tau)$ $=\frac{1}{\mathbb{E}[a(X)\mid W_{0}=1]}\sum_{k=1}^{K}a(x_{k})\tau(x_{k})\mathbb{P}(X=x_{k}\mid W_{0}=1)\geq 0.$
The inequality holds because $a(x_{k})$ and $\tau(x_{k})$ are nonnegative for all $k\in\{1,\ldots,K\}$. This last inequality is reversed if we instead assume that $\tau\leq\textbf{0}_{K}$. Thus, $\mu(a,\tau_{0})$ is weakly causal. ∎
Appendix G Proofs for Appendix C
We use the following lemma in the proof of Theorem C.1.
Lemma G.1.
Let $\theta=(\theta(1),\ldots,\theta(K))\in\mathbb{R}^{K}$ and define the mapping $\phi:\mathbb{R}^{K}\to\mathbb{R}$ by $\phi(\theta)=\max_{j\in\{1,\ldots,K\}}\theta(j)$. Then, $\phi$ is Hadamard directionally differentiable for all $\theta\in\mathbb{R}^{K}$ tangentially to $\mathbb{R}^{K}$ with directional derivative at $\theta$ in direction $h\in\mathbb{R}^{K}$
$\phi_{\theta}^{\prime}(h)$ $=\max_{j\in\operatorname*{arg\,max}_{k\in\{1,\ldots,K\}}\theta(k)}h(j).$
Proof of Lemma G.1.
Let $h_{n}\rightarrow h\in\mathbb{R}^{K}$ and $t_{n}\searrow 0$ as $n\rightarrow\infty$. Then,
$t_{n}^{-1}(\phi(\theta+t_{n}h_{n})-\phi(\theta))$ $=t_{n}^{-1}\left(\max_{k\in\{1,\ldots,K\}}(\theta(k)+t_{n}h_{n}(k))-\max_{k\in\{1,\ldots,K\}}\theta(k)\right).$
Let $\Theta_{\max}=\{j\in\{1,\ldots,K\}:\theta(j)=\max_{k\in\{1,\ldots,K\}}\theta(k)\}$ and let $j_{\max}$ be an element of $\Theta_{\max}$. Then, $\max_{k\in\{1,\ldots,K\}}\theta(k)=\theta(j_{\max})$ and thus
$\frac{\phi(\theta+t_{n}h_{n})-\phi(\theta)}{t_{n}}$ $=\max\left\{\frac{\theta(1)-\theta(j_{\max})}{t_{n}}+h_{n}(1),\ldots,\frac{\theta(K)-\theta(j_{\max})}{t_{n}}+h_{n}(K)\right\}.$
For each $j\in\Theta_{\max}$, $(\theta(j)-\theta(j_{\max}))/t_{n}+h_{n}(j)=h_{n}(j)\rightarrow h(j)$. For each $j\notin\Theta_{\max}$, $(\theta(j)-\theta(j_{\max}))/t_{n}\rightarrow-\infty$ since $\theta(j)-\theta(j_{\max})<0$ and $t_{n}\searrow 0$. Therefore, by continuity of the maximum operator in its arguments, $t_{n}^{-1}(\phi(\theta+t_{n}h_{n})-\phi(\theta))\rightarrow\max_{j\in\Theta_{\max}}h(j)$. ∎
Proof of Theorem C.1.
We begin by showing the consistency of $\widehat{\overline{P}}$ for $\overline{P}$.
Part 1: Consistency
The estimator $\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})$ is consistent for $\mathbb{E}[a(X)w_{0}(X)]$ since its components are assumed consistent by Assumption C.1, and by the continuous mapping theorem. The consistency of $\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})$ for $\mathbb{E}[w_{0}(X)]$ is similarly established.
We now consider the maximum term in the denominator. We can write
$\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})$ $=\max_{x:\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x)>0,\widehat{w}_{0}(x)>c_{n}}\widehat{a}(x).$
Let $\mathcal{X}^{+}\coloneqq\text{supp}(X\mid W_{0}=1)=\{x\in\text{supp}(X):w_{0}(x)>0\}$ and let $\widehat{\mathcal{X}}^{+}=\{x:\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x)>0,\widehat{w}_{0}(x)>c_{n}\}$. We first show that $\mathbb{P}(\widehat{\mathcal{X}}^{+}=\mathcal{X}^{+})\rightarrow 1$ as $n\rightarrow\infty$. To see this, first consider $x_{j}\in\mathcal{X}^{+}$. Then,
$\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})$ =($=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0\right\}\cap\left\{\widehat{w}_{0}(x_{j})>c_{n}\right\}\right)$
$\geq\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0\right)+\mathbb{P}\left(\widehat{w}_{0}(x_{j})>c_{n}\right)-1,$
following an application of Bonferroni’s inequality.